OpenStudy (anonymous):

Determine whether the sequence converges or diverges. If it converges, give the limit. 60, -10, 5/3, -5/18, ...

OpenStudy (anonymous):

I think it converges cause it keeps getting smaller

OpenStudy (anonymous):

If we ignore the negative terms for the moment, can you see that every term is getting smaller and smaller by a factor of \(\dfrac{1}{6}\)? Meaning, \[60\times\frac{1}{6}=10\\ 10\times\frac{1}{6}=\frac{10}{6}=\frac{5}{3}\\ \frac{5}{3}\times\frac{1}{6}=\frac{5}{18}\] and so on.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Okay, so if you keeping scaling a number down like this what value do you approach?

OpenStudy (anonymous):

0?

OpenStudy (anonymous):

Exactly right! If we accounted for the negative, we'd have a factor of \(-\dfrac{1}{6}\), but that won't change that conclusion. What we've determined is that we can model this sequence with the formula \[a_n=60\left(-\frac{1}{6}\right)^n\] which is a geometric sequence. If you can determine that the ratio makes the next terms approach 0, then the sequence converges. If you want a more formal approach to proving the limit exists, you can show the sequence is bounded and "monotonic", or that the sequence is Cauchy, but something tells me that's a bit beyond the scope of the question.

OpenStudy (anonymous):

so it converges and the limit is 0

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Awesome, thanks!!!! :D

OpenStudy (anonymous):

yw!